Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
mwas, tomelleri - 13:45 Tuesday 18 November 2025 (68183) Print this report
OMC slow shutter driver replacement

The OMC slow shutter did not want to open this morning. 

Figure 1. Looking at the OMC temperature band passed at 1.7kHz (fast RMS) and band passed at 666Hz (slow RMS), the electrical signal to move the translation stage of the shutter is no longer being sent. Normally there is a cross talk between the sent voltage to make the motor steps, and the temperature sensor. The last successful commands to move the shutter where on November 17 before 6 UTC, before the electrical problems at the site. Afterwards there have been many commands sent on Nov 17 in the afternoon by the automation, which did not have an effect. 

First the cable of the OMC slow shutter was moved from channel 2 axis 1 to channel 4 axis 1 around 9:15 UTC. This did not change the situation, there was still no cross talk visible on the OMC temperature. Trying to rotate the slow shutter of the Faraday isolator showed that the driver does send commands on channel 4 axis 2. Then the cables of the Faraday isolator and the OMC slow shutter were swapped around 9:45 UTC, so that the OMC slow shutter is now on channel 4 axis 2. After that there started to be electrical signals sent to the OMC slow shutter visible in the OMC temperature, but the shutter is still not moving and not letting light go through to the OMC.  

We then swapped back the OMC shutter and FI shutter cables around again. And looking at the position of the FI shutter it was also not moving, even though the electrical signals were being sent to it.

Channel 2 axis 1 and Channel 4 axis 1 appear to not work on the Agilis driver, FT4ZNKQD. Looking back into the logbook that was already known. Next time we should put some tape on the broken channels so it is more evident. The whole driver was also not moving the rotators, even though the cross talk of the driving signal was visible. Either the driver is fully not functional, or it needed more actions to recover, like turning it off and on. We haven't tested which one of the two hypothesis is true.

The driver was replaced by a spare, FT6G3NUZ. This has solved the problem, both the FI shutter and the OMC shutter started to move again. The cabling has not been updated, so now the OMC slow shutter is connected to channel 4 axis 1, instead of channel 2 axis 1. The buttons in VPM have been updated, as well as DET_MAIN.py and ITF_LOCK.py to use the new channel.

After that I have also check that the OMC lock works, and update the demodulation phase of B1 PD3 and B1 by 0.5 radian, so that all of the signal is on the I quadrature, instead of half on I and half on Q.

2025-11-17-05h24m39-UTC>INFO...-Received move_rel message from DET_MAIN
2025-11-17-05h24m39-UTC>INFO...-Parsed message*: [2, 1, 45000, 50]
2025-11-17-05h24m39-UTC>INFO...-setting channel 2 axis 1 to 45000 at speed 50
2025-11-17-05h24m39-UTC>INFO...-writing position -103695199
2025-11-17-05h56m16-UTC>INFO...-Received move_rel message from DET_MAIN
2025-11-17-05h56m16-UTC>INFO...-Parsed message*: [2, 1, -60000, 50]
2025-11-17-05h56m16-UTC>INFO...-setting channel 2 axis 1 to -60000 at speed 50

2025-11-17-13h07m18-UTC>INFO...-Received move_rel message from DET_MAIN
2025-11-17-13h07m18-UTC>INFO...-Parsed message*: [2, 1, -60000, 50]
2025-11-17-13h07m18-UTC>INFO...-setting channel 2 axis 1 to -60000 at speed 50
2025-11-17-13h07m18-UTC>INFO...-writing position -103815199
2025-11-17-13h09m15-UTC>INFO...-Received move_rel message from DET_MAIN
2025-11-17-13h09m15-UTC>INFO...-Parsed message*: [2, 1, -60000, 50]
2025-11-17-13h09m15-UTC>INFO...-setting channel 2 axis 1 to -60000 at speed 50
2025-11-17-13h09m15-UTC>INFO...-writing position -103875199
2025-11-17-13h11m11-UTC>INFO...-Received move_rel message from DET_MAIN
2025-11-17-13h11m11-UTC>INFO...-Parsed message*: [2, 1, -60000, 50]
2025-11-17-13h11m11-UTC>INFO...-setting channel 2 axis 1 to -60000 at speed 50
2025-11-17-13h11m11-UTC>INFO...-writing position -103935199

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×